OpenStudy (babyslapmafro):

Please help me solve the following.... f ' (x)=(sqrt x) '+ (4/x) '

OpenStudy (bahrom7893):

what are we supposed to find?

OpenStudy (babyslapmafro):

f-prime of x

OpenStudy (bahrom7893):

oh I see. so they just want you to find the derivative

OpenStudy (babyslapmafro):

yes, sorry that was not clear enough

OpenStudy (bahrom7893):

So.. f'(x) = (x^(1/2)+(4x^-1))'

OpenStudy (bahrom7893):

= (1/2)x^(-1/2) - 4x^(-2)

OpenStudy (bahrom7893):

or f'(x) = 1/(2sqrt(x)) - 4/(x^2)

OpenStudy (babyslapmafro):

nice picture-swap lol

OpenStudy (babyslapmafro):

thank you for the help, my algebra is pathetic so i guess i must review what you did in greater detail

OpenStudy (bahrom7893):

sqrt(x) = x^(1/2) 1/x^a = x^(-a)

OpenStudy (babyslapmafro):

thanks